Consider a cube of material of sides 1 cm with a resistance of 0.003 O between its opposite faces. What is the resistance if the same volume of the material is drawn into a length of 50cm with a uniform cross – section?

Grade:10

2 Answers

Vikas TU
7 years ago
For 1 cm cube,
Resistance is=>
R = pl/A
0.003 = p*1/(1*1)..............(1)
For 50 cm cube,
R’ = p*50/(50*50)....................(2)
Dividing both the eqns. we get,
R’ = 0.003/50 => 0.6 x 10^-4 ohm.
shobitha
7 years ago
R = pl/A
0.003 = p*1 cm/(V)
0.003= p/(V)..............(1)
For 50 cm cube,for same volume of the material divide volume(v) by 50
R’ = p*50/(v/50)
R’ = p*2500/(v)....................(2)
solving (1) and (2)
we get R'=7.5 ohm